OpenStudy (anonymous):

can someone help me please! use the elimination method to solve the system of equations. a+7b=16 a+3b=4 a. there is one solution, it is____ b. there are infinitely many solutions. c. there is no solution.

OpenStudy (anonymous):

Hi, a+7b=16 a+3b=4 ||First subtract 2nd EQ from te 1st EQ to eliminate a 4b = 12 b = 3 The solve for a using one of the EQ above , a = -5 Ordered pair(-5,3) is the solution for this system of equations

OpenStudy (mathstudent55):

Subtract equation 2 froim equation 1: a + 7b = 16 a + 3b = 4 -------------- 4b = 12 Then b = 3 Now substitute 3 in for b in original eqation 1: a + 7(3) = 16 a + 21 = 16 a = -5

OpenStudy (anonymous):

so the solution is a= - 5?

OpenStudy (mathstudent55):

The solution is : a = -5, b = 3 You need to state what both a and b are.

OpenStudy (anonymous):

hey mathstudent55 could you help me with my question please?

OpenStudy (anonymous):

okay thank you very much

OpenStudy (anonymous):

so would the correct answer that I type be, (-5,3)?

OpenStudy (mathstudent55):

yes

OpenStudy (anonymous):

okay thank you for your help.